A golf ball has a mass 40g speed 45m\s calculate its wavelength

Grade:11

1 Answers

Vikas TU
7 years ago
From Hiesenberg’s Uncertainity princple,
p*x > = h/4pi
(40/1000)*45*x = h/4pi.
put h and pi value and get x.
that would be the error/uncertainity in position x .i.e. wavelength.